# Pyramid Vector Quantization for LLMs

### v1

@@ 3.1 Why use PVQ for LLM quantization?
Figure 3: Signal-to-noise-ratio of weight and activation quantization methods on a Gaussian source. PVQ achieves similar high signal-to-noise as state-of-the-art E8 weight-only vector quantization. As PVQ is search-free and codebook-free, it can practically be used to quantize both weight and activations.
3.2 Signal-to-noise on ideal Gaussian source
To get an idea of how effective different quantization algorithms are, we start by comparing their performance on ideal Gaussian source data. We measure the signal-to-noise ratio by averaging the mean squared error between true and quantized signal over 1000 samples in Figure 3. We find that PVQ achieves high signal-to-noise ratios on Gaussian data equivalent to E8, a state-of-the-art quantization method that exploits geometric structure. PVQ achieves especially superior signal-to-noise ratios for activation quantization, in which case we compare to RTN without search as a baseline that is suitable to activation quantization also.
@@ 3.3 Pyramid vector quantization for LLMs
In PVQ, the trade-off between performance and the effective bits per weight (BPW) can be controlled through the groupsize $D$, shape bits $b_{\text{shape}}$ and gain bits $b_{\text{gain}}$:
$\text{BPW}=b_{\text{shape}}+b_{\text{gain}}/G$ (7)
The average shape bits $b_{\text{shape}}$ are not required to be integer, and can be a fractional (e.g. $b_{\text{shape}}=3.5$).
@@
To determine the number of pulses $K$ such that the number of bits required to encode PVQ vectors remains within the desired maximum, we find the largest $K$ by increasing it such that it still satisfies $\left\lceil\log_{2}(N(D,K))/D\right\rceil<b_{\text{shape}}$. Here, $N(D,K)$ is computed using the recursive algorithm outlined in Section 2.4. In practice, the number $N(D,K)$ can exceed regular integer types and may require arbitrary precision integers.
Step 3. Quantizing the shape.
[…]
Step 4. Computing the gain.
For each quantized row vector $\widehat{{\bm{w}}}\in\mathbb{R}^{D}$ in a group, we can find the optimal scale parameter $s\in\mathbb{R}$ that minimizes the squared error to the original weight matrix ${\bm{w}}$ in closed-form $s=\widehat{{\bm{w}}}^{T}\widehat{{\bm{w}}}/\widehat{{\bm{w}}}^{T}{\bm{w}}$. Repeating this operation across features and groups results in a gain matrix ${\bm{S}}\in\mathbb{R}^{N\times G}$. We refer to the quantized $\widehat{{\bm{W}}}$ as the ‘shape’ and the scales ${\bm{S}}$ as the ‘gain’.
Step 5. Quantizing the gain (optional)
For small groupsizes, we may also consider quantizing the gain using the approach discussed in Section 3.4. For each row ${\bm{s}}\in\mathbb{R}^{G}$ in ${\bm{S}}$, this entails quantizing normalized elements $\lfloor\text{CDF}(s_{i}^{2}/||{\bm{s}}||_{2}^{2})\cdot 2^{b}\rfloor$ using the CDF of the $\text{Beta}(D/2,D(G-1)/2)$ distribution. The normalizing constant $||{\bm{s}}||_{2}^{2}$ needs to be stored for dequantization, but because it is shared across groups its contribution to the total bits per weight $<0.01$ is negligibly small.
@@
After quantizing each group of weights $\widehat{{\bm{W}}}_{g}$, we can perform a partial update to the remaining weights ${\bm{W}}_{>g}:=\begin{bmatrix}{\bm{W}}_{g+1}&{\bm{W}}_{g+2}\ldots{\bm{W}}_{G}\end{bmatrix}\in\mathbb{R}^{N\times(G-g)D}$ to ‘compensate’ the error of the quantization ${\bm{W}}_{g}$ that best minimizes our proxy loss $L$ of Equation 1. To compute the update, we use the transpose of the inverse Cholesky decomposition ${\bm{L}}^{-T}$ of the Hessian ${\bm{H}}={\bm{L}}{\bm{L}}^{T}$.
${\bm{W}}_{>g}\leftarrow({\bm{W}}_{g}-{\bm{s}}_{g}^{T}..\text{(think about most elegant notation here)}...\widehat{{\bm{W}}}_{g})...{\bm{L}}^{-T}....$ (8)
The correction was originally proposed in the context of scalar quantization in GPTQ (Frantar et al., 2022), but is also applicable in pyramid vector quantization, as we demonstrate.
Step 7. Encoding the shape.
To encode a quantized weight $\widehat{{\bm{W}}}\in\mathbb{R}^{N\times GD}$ (e.g. 32 bit float for $\mathbb{R}$) into the actual low bit integer representation $\overline{{\bm{W}}}\in\mathbb{N}^{N\times G}$, sometimes referred to as the ‘code’, we can use the PVQ encoding algorithm described in Algorithm 2. In practice, the amount of bits required to represent a code may grow beyond the maximum integer precision of many languages (e.g. when $\lceil\log_{2}(N(D,K))\rceil>128$ bits, it can not be represented in a 128 bit integer). We implemented a hardware kernels for encoding and decoding using custom arithmetic precision arithmetic in CUDA/C++.
3.4 Gain quantization
Normalized gains follow the Beta distribution.
Smaller groupsizes lead to a lower quantization error, but lead to an increase in the number of bits required to store gain parameters. This is because the gain ${\bm{S}}\in\mathbb{R}^{N\times G}$ grows linearly with the number of groups $G$, and therefore inversely proportional to the chosen groupsize $D$. To overcome this issue, we propose a theoretically and empirically motivated scheme to quantize gain parameters. In Theorem 3.1, we note that row vectors of normalised gains follow a Beta distribution $\text{Beta}(\frac{D}{2},\frac{D(G-1)}{2})$ of which the coefficients depend on the groupsize $D$ and the number of groups $G$. We empirically confirm that our theory meets practice, as we find that the Beta distribution matches normalized weight distributions of pretrained LLM models, after performing the rotation described in Section 2.3, as shown for selection of layers in a pretrained LLM in Figure 4. To exploit the observation that gains are Beta distributed, we propose to use the quantiles of this distribution to quantize gains.
Figure 4: Effect of different shape and gain bits on effective PVQ quantization grid. The distribution automatically matches the theoretical Beta weight distribution.
@@ 3.4 Gain quantization
Let ${\bm{w}}\in\mathbb{R}^{GD}$ be a normally distributed vector that can be grouped in $G$ equally sized vectors ${\bm{w}}=\begin{bmatrix}{\bm{v}}_{1}&{\bm{v}}_{2}&\cdots&{\bm{v}}_{G}\end{bmatrix}$ where each of the vectors ${\bm{v}}_{g}$ has the same dimensionality equal to the groupsize ${\bm{v}}_{g}\in\mathbb{R}^{D}$. Then the normalised radius (the ‘gain’) of each group $s_{h}={\bm{v}}_{g}^{T}{\bm{v}}_{g}/||{\bm{w}}||_{2}^{2}$ follows the $s_{g}\sim\text{Beta}(\frac{D}{2},\frac{G(D-1)}{2})$ distribution. (Proof in Section A.1)
Figure 5: Theoretical Beta distribution of Theorem 3.1 closely match gains of rotated weights in trained LLMs. Here demonstrated for empirical weight distributions of a pretrained Llama-v3 8B.
@@
Quantizing gains using Beta quantiles.
The observation that normalized gains are Beta distributed is important, as it suggests that gains can effectively be quantized with the quantiles of this Beta distribution, without having to introduce additional hyper-parameters. Effectively, $b$ bit quantization using quantiles amounts to quantizing on a regular uniform grid of $2^{b}$ points after first transforming elements through a change-of-variables given by the CDF of the Beta distribution and dequantizing using the inverse CDF (see Figure 6):
@@ 4 Results
4.1 Weight-only quantization (without gain quantization)
Since most prior work focuses on weight-only quantization, we begin by comparing the performance of weight-only PVQ quantization with common weight-only quantization baselines. Following prior work (Frantar et al., 2022; Ashkboos et al., 2024), we measure test perplexity (PPL) and average accuracy on a range of zero-shot downstream tasks after quantizing common open-source LLM models of Phi, Mixtral and Llama families. We compare groupsizes in [16, 32, 64, 128, 256], keep the gain in 16 bit, and use shape bits in [3, 3.5, 4, 4.5, 5, 5.5, 6, 7, 8] for PVQ and [3, 4, 5, 6, 7, 8] for other methods. PVQ is the only quantization method that supports non-integer shape bits.
Figure 7: Weight only quantization. Test perplexities (PPL) after quantizing with different methods at group size settings (within connected set) and various shape bits (between connected sets).
@@ 4.1 Weight-only quantization (without gain quantization)
4.2 Weight-only quantization (shape and gain quantization)
A benefit of PVQ is that the number of bits for the shape and bits for the gains can be chosen flexibly, even to not-integer ratios. To evaluate the effect of different gain bits, we repeat the same experiment as before, but rather than varying the groupsize, we fix the groupsize to 16 and vary the amount of bits used for the gain.
Figure 8: Shape and gain quantization. Test perplexities (PPL) and test accuracies (Avg. Acc ) at different bits for gains (within connected sets) and various shape bits (between sets).
@@ 4.2 Weight-only quantization (shape and gain quantization)
PVQ [2.5 bit shapes, 16 bit gains] 16 3.50 7.52 0.67 4.42 0.76 8.04 0.68 5.92 0.74
PVQ [3 bit shapes, 4 bit gains] 16 3.25 6.85 0.69 4.22 0.76 7.14 0.70 4.51 0.78
Table 2: Quantizing shape and gain. We compare post-training perplexity (PPL) and average zero-shot performance (Avg. Acc). PVQ yields the highest performance after quantization. Details and additional results in Section E.2.
@@ 4.3 Weights and Activation quantization
As PVQ is codebook-free and search-free, it can be applied to not only the weights but also the activations during inference. In Figure 13 we evaluate the final test perplexity for different settings of effective bits per weight and bits per activations. In Table 3, we compare the resulting test perplexities to other methods for quantizing weight and activation. We also include quantization methods that are not codebook-free, by relying on round-to-nearest (RTN) to quantize activations in real-time during inference. We find that PVQ obtains state-of-the-art performance across the considered LLM architectures, and that this holds generally across different settings of bit rates.
@@ 4.4 Downstream zero-shot tasks
In Table 4, we provide results on downstream zero-shot tasks split out per task. We report weight-only PVQ with both shape and gain quantization on a Llama-3-8B model. Additional results including other LLM models can be found in Section E.4.
@@
Original 16.00 6.01 0.81 0.73 0.78 0.79 0.57 0.65 0.72
RTN 4.00 8.39 0.77 0.68 0.74 0.76 0.53 0.55 0.67
GPTQ 16 4.00 6.82 0.79 0.71 0.74 0.79 0.57 0.63 0.70
QuaRot 16 4.00 6.66 0.78 0.72 0.73 0.78 0.55 0.64 0.70
PVQ [3.0 bit shapes, 4 bit gains] 16 3.25 6.85 0.79 0.72 0.74 0.77 0.54 0.60 0.69
PVQ [3.5 bit shapes, 6 bit gains] 16 3.88 6.33 0.81 0.72 0.76 0.80 0.57 0.64 0.72
@@ 4.5 Optimizing PVQ for CUDA
The recurrent formulation for the size table ($N\left[D,K\right]$) quickly becomes prohibitive as its naive implementation requires $O(D\cdot K\cdot W)$ serial operations, each requiring overlapping reads and writes. Instead, we reformulate the algorithm to perform $K$ operations in parallel using a CUDA optimized Hillis-Steele type scan reduction (Hillis & Steele, 1986) which accounts for the $\log(K)$ factor in Table 5, computing each row in parallel (Section C.1). This has the further advantage of only having to read from global memory for synchronization following (Xiao & Feng, 2010), otherwise the threads keep the values in registry and exclusively writes to global memory.
@@
Encoding / Decoding $O(B\cdot D\cdot W)$ $O(B\cdot D\cdot W)$ reads, $O(B\cdot D)$ ordered writes
Size table generation $O(K\cdot\log(K)\cdot D\cdot W)$ $O(K\cdot D\cdot W)$ ordered writes
Table 5: The time and I/O complexity on the HBM for our CUDA optimized PVQ kernels, where $B$ is the batch size and $W$ the number of words comprising each integer. All functions have $O(W)$ space complexity. We highlight ordered I/O operations as this more efficient use of the memory bandwidth as they can be coalesced into fewer operations as opposed to random ones.

### v2

@@ 3.1 Practical advantages of PVQ
Figure 3: Signal-to-quantization-noise-ratio (QSNR) of quantization methods on standard Gaussian source. PVQ achieves high QSNR close to E8, which uses an optimal packing on a uniform source. PVQ uses an implicit codebook and is search-free, thereby amenable to quantization of weight and activations.
3.2 Signal-to-quantization-noise on ideal Gaussian source
To assess the theoretical effectivity between different quantization algorithms are, we start by comparing empirical estimates of their performance on an idealized standard Gaussian source, zero mean unit variance. We measure the signal-to-quantization-noise ratio QSNR by averaging the mean squared error between true and quantized signal over 1000 samples in Figure 3. We compare PVQ with E8, a method that has optimal packing of uniformly distributed weights in $D=8$ dimensions only (Tseng et al., 2024a) and naive rounding-to-nearest (RTN) scalar quantization. We find that PVQ achieves QSNR ratios close to the optimal E8 method. Further, PVQ can also be applied to activation quantization, in which case we compare to RTN without search as a baseline that is suitable to quantization of both weights and activations.
@@ 3.3 Pyramid vector quantization for LLMs
In PVQ, the trade-off between performance and effective bits per weight (BPW) is controllable through the groupsize $D$, which must divide the number of columns so that ${\bm{W}}\in\mathbb{R}^{N\times DG}$, direction bits $b_{\text{direction}}\in\mathbb{N}$ and amplitude bits $b_{\text{amplitude}}\in\mathbb{N}$:
$\text{BPW}=b_{\text{direction}}/D+b_{\text{amplitude}}/G$ (7)
As the number of direction bits per weight $b_{\text{direction}}/D$, and amplitude bits per weight $b_{\text{amplitude}}/G$ are fractions, it is easy to choose a non-integer number of bits per weight in PVQ.
@@
To determine the number of pulses $K$ such that the number of bits required to encode PVQ vectors remains within the desired maximum, we find the largest $K$ by increasing it such that it still satisfies $\left\lceil\log_{2}(N(D,K))/D\right\rceil<b_{\text{direction}}$. Here, $N(D,K)$ is computed using the recursive algorithm outlined in Section 2.4. In practice, the number $N(D,K)$ can exceed regular integer types and may require arbitrary precision integers.
Step 3. Quantizing the direction.
We quantize all LLM weight matrices making up key, query, value and fully-connected components. We write the normalized weight matrix ${\bm{W}}=\begin{bmatrix}{\bm{W}}_{1}&{\bm{W}}_{2}&\ldots&{\bm{W}}_{G}\end{bmatrix}\in\mathbb{R}^{N\times GD}$, with $N$ features, $G$ groups and a groupsize of $D$. Given our choice of $K$, we quantize groups ${\bm{W}}_{g}\in\mathbb{R}^{N\times D}$ using the quantization procedure of Section 2.4 yielding a quantized direction matrix $\widehat{{\bm{W}}}_{g}$ on the pyramid ${\mathcal{P}}_{D,K}$, ie. elements are rounded integers and absolute values of rows sum to 1. This operation can be parallelized over features and groups.
Step 4. Computing the amplitude.
For each quantized row vector $\widehat{{\bm{w}}}\in\mathbb{R}^{D}$ in a group, we can find an optimal rescaling by $s\in\mathbb{R}$ that minimizes the Euclidean distance to the original weight ${\bm{w}}$ in closed-form $s=\widehat{{\bm{w}}}^{T}\widehat{{\bm{w}}}/\widehat{{\bm{w}}}^{T}{\bm{w}}$. Repeating this for all features and groups yields an amplitude matrix ${\bm{S}}\in\mathbb{R}^{N\times G}$. We refer to quantized $\widehat{{\bm{W}}}$ as the ‘direction’ and the scales ${\bm{S}}$ as the ‘amplitude’.
Step 5. Quantizing the amplitude (optional)
For small groupsizes, we can optionally quantize amplitudes as described in Section 3.4. For each row ${\bm{s}}\in\mathbb{R}^{G}$ in ${\bm{S}}$, this entails quantizing normalized elements $\lfloor\text{CDF}(s_{i}^{2}/||{\bm{s}}||_{2}^{2})\cdot 2^{b}\rfloor$ using the CDF of the $\text{Beta}(D/2,D(G-1)/2)$ distribution. The normalizing constant $||{\bm{s}}||_{2}^{2}$ needs to be stored for dequantization, but because it is shared across groups its contribution to the total $<0.01$ bits per weight is negligibly small.
@@
For each quantized group of weights indexed by the quantized columns ${\bm{s}}_{g}\widehat{{\bm{W}}}_{g}\in\mathbb{R}^{N\times D}$, we can update the other remaining columns ${\bm{W}}_{\neg g}\in\mathbb{R}^{N\times(G-1)D}$ to compensate the quantization error that minimizes the proxy loss $L$ of Equation 1:
${\bm{W}}_{\neg g}\leftarrow{\bm{W}}_{\neg g}-{\bm{H}}^{-1}_{\neg g,\neg g}{\bm{H}}_{\neg g,g}({\bm{W}}_{g}-{\bm{s}}_{g}\widehat{{\bm{W}}}_{g})$ (8)
where ${\bm{H}}_{\neg g,\neg g}$ is a square submatrix of the inverse Hessian with rows and columns corresponding to remaining weights and ${\bm{H}}^{-1}$ is a rectangular submatrix of the inverse Hessian with rows and columns that correspond to quantized and remaining weights respectively. We can avoid inverting the full Hessian for every update by working with its Cholesky decomposition ${\bm{H}}={\bm{L}}{\bm{L}}^{T}$, as detailed in GPTQ (Frantar et al., 2022) which proposed the update in the context of scalar quantization.
Step 7. Encoding the direction.
To encode a quantized weight $\widehat{{\bm{W}}}\in\mathbb{R}^{N\times GD}$ (e.g. 32 bit float for $\mathbb{R}$) into the actual low bit integer representation $\overline{{\bm{W}}}\in\mathbb{N}^{N\times G}$, sometimes referred to as the ‘code’, we can use the PVQ encoding algorithm described in Algorithm 2. This is problematic, since the number of bits required to store a vector may exceed the bit width of integer types in many languages (e.g. when $\lceil\log_{2}(N(D,K))\rceil>128$ bits, it can not be represented in a 128 bit integer). To overcome this, we implemented arbitrary precision arithmetic in CUDA/C++ to support arbitrary bit width integer types to allow encoding and decoding kernels.
3.4 amplitude quantization
Normalized amplitudes follow the Beta distribution.
Smaller groupsizes lead to a lower quantization error, but lead to an increase in the number of bits required to store amplitude parameters. This is because the amplitude ${\bm{S}}\in\mathbb{R}^{N\times G}$ grows linearly with the number of groups $G$, and therefore inversely proportional to the chosen groupsize $D$. To overcome this issue, we propose a theoretically and empirically motivated scheme to quantize amplitude parameters. In Theorem 3.1, we note that row vectors of normalized amplitudes follow a Beta distribution $\text{Beta}(\frac{D}{2},\frac{D(G-1)}{2})$ of which the coefficients depend on the groupsize $D$ and the number of groups $G$. We empirically confirm that our theory meets practice, as we find that the Beta distribution matches normalized weight distributions of pretrained LLM models, after performing the rotation described in Section 2.3, as shown for selection of layers in a pretrained LLM in Figure 4. To exploit the observation that amplitudes are Beta distributed, we propose to use the quantiles of this distribution to quantize amplitudes.
Figure 4: Effect of different direction and amplitude bits on effective PVQ quantization grid. The distribution automatically matches the theoretical Beta weight distribution.
@@
Let ${\bm{w}}\in\mathbb{R}^{GD}$ be a normally distributed vector that can be grouped in $G$ equally sized vectors ${\bm{w}}=\begin{bmatrix}{\bm{v}}_{1}&{\bm{v}}_{2}&\cdots&{\bm{v}}_{G}\end{bmatrix}$ where each of the vectors ${\bm{v}}_{g}$ has the same dimensionality equal to the groupsize ${\bm{v}}_{g}\in\mathbb{R}^{D}$. Then the normalized radius (the ‘amplitude’) of each group $s_{g}={\bm{v}}_{g}^{T}{\bm{v}}_{g}/||{\bm{w}}||_{2}^{2}$ follows the $s_{g}\sim\text{Beta}(\frac{D}{2},\frac{G(D-1)}{2})$ distribution. (Proof in Section A.1)
Figure 5: Theoretical Beta distribution of Theorem 3.1 closely match amplitudes of rotated weights in trained LLMs, here demonstrated for empirical weight distributions of a pretrained Llama-v3 8B.
@@
Quantizing amplitudes using Beta quantiles.
The observation that normalized amplitudes are Beta distributed is important, as it suggests that we can quantize amplitudes efficiently by mapping centers of linearly spaced regions through the quantile function of a Beta distribution, without introducing additional hyper-parameters. To obtain a $b$ bit quantizer, we take a regular uniform grid of $2^{b}$ points after first transforming elements through a change-of-variables given by the CDF of the Beta distribution and dequantizing using the inverse CDF (see Figure 6):
@@ 4 Results
4.1 Weight-only quantization (without amplitude quantization)
Since most prior work focuses on weight-only quantization, we begin by comparing the performance of weight-only PVQ quantization with common weight-only quantization baselines. Following prior work (Frantar et al., 2022; Ashkboos et al., 2024), we measure test perplexity (PPL) and average accuracy on a range of zero-shot downstream tasks after quantizing common open-source LLM models of Phi, Mixtral and Llama families. We compare groupsizes in [16, 32, 64, 128, 256], keep the amplitude in 16 bit, and use direction bits in [3, 3.5, 4, 4.5, 5, 5.5, 6, 7, 8] for PVQ and [3, 4, 5, 6, 7, 8] for other methods. Unlike most scalar quantization methods, PVQ more easily supports non-integer number of direction bits as we can choose an integer number of bits per codeword that is not divisible by the number of groups $G$ – for instance, with groupsize $D=16$ and 40 bits per group results in $40/16=2.5$ direction bits).
Figure 7: Weight only quantization. Test perplexities (PPL) after quantizing with different methods at group size settings (within connected set) and various direction bits (between connected sets).
@@ 4.1 Weight-only quantization (without amplitude quantization)
4.2 Weight-only quantization (direction and amplitude quantization)
A benefit of PVQ is that the number of bits for the direction and bits for the amplitudes can be chosen flexibly, even to not-integer ratios. To evaluate the effect of different amplitude bits, we repeat the same experiment as before, but rather than varying the groupsize, we fix the groupsize to 16 and vary the amount of bits used for the amplitude.
Figure 8: direction and amplitude quantization. Test perplexities (PPL) and test accuracies (Avg. Acc ) at different bits for amplitudes (within connected sets) and various direction bits (between sets).
@@ 4.2 Weight-only quantization (direction and amplitude quantization)
PVQ [2.5 bit directions, 16 bit amplitudes] 16 3.50 7.52 0.67 4.42 0.76 8.04 0.68 5.92 0.74
PVQ [3 bit directions, 4 bit amplitudes] 16 3.25 6.85 0.69 4.22 0.76 7.14 0.70 4.51 0.78
Table 2: Quantizing direction and amplitude. We compare post-training perplexity (PPL) and average zero-shot performance (Avg. Acc). PVQ yields the highest performance after quantization. Details and additional results in Section E.2.
@@ 4.3 Weights and Activation quantization
As PVQ use an implicit codebook and is search-free, it can be applied to not only the weights but also the activations during inference and reduce computational requirements of the forward-pass. In Figure 13 we evaluate the final test perplexity for different settings of effective bits per weight and bits per activations. In Table 3, we compare the resulting test perplexities to other methods for quantizing weight and activation. We also compare to naive round-to-nearest (RTN) without search, which is also amenable to quantization of weights and activations but has much lower signal to noise. We find that PVQ obtains state-of-the-art performance across the considered LLM architectures, and that this holds generally across different settings of bit rates.
@@ 4.4 Downstream zero-shot tasks
In Table 4, we provide results on downstream zero-shot tasks split out per task. We report weight-only PVQ with both direction and amplitude quantization on a Llama-3-8B model. Additional results including other LLM models can be found in Section E.4.
@@
Original 16.000 6.13 0.81 0.73 0.79 0.78 0.53 0.76 0.73
RTN 3.125 29.41 0.64 0.55 0.42 0.41 0.25 0.22 0.41
GPTQ 128 3.125 17.77 0.63 0.59 0.35 0.43 0.26 0.17 0.40
QuaRot 128 3.125 7.62 0.77 0.71 0.73 0.75 0.46 0.71 0.69
PVQ [3 bit directions, 16 bit amplitudes] 128 3.125 7.01 0.80 0.73 0.76 0.78 0.50 0.75 0.72
@@ 4.5 Optimizing PVQ for CUDA
The recurrent formulation for the size table ($N\left[D,K\right]$) quickly becomes prohibitive as its naive implementation requires $O(D\cdot K\cdot G)$ serial operations, each requiring overlapping reads and writes. Instead, we reformulate the algorithm to perform $K$ operations in parallel using a CUDA optimized Hillis-Steele type scan reduction (Hillis & Steele, 1986) which accounts for the $\log(K)$ factor in Table 5, computing each row in parallel (Section C.2). This has the further advantage of only having to read from global memory for synchronization following (Xiao & Feng, 2010), otherwise the threads keep the values in registry and exclusively writes to global memory.
@@
Encoding / Decoding $O(B\cdot D\cdot G)$ $O(B\cdot D\cdot G)$ reads, $O(B\cdot D)$ ordered writes
Size table generation $O(K\cdot\log(K)\cdot D\cdot G)$ $O(K\cdot D\cdot G)$ ordered writes
Table 5: The time and I/O complexity on the HBM for our CUDA optimized PVQ kernels, where $B$ is the batch size and $G$ the number of words comprising each integer. All functions have $O(G)$ space complexity. We highlight ordered I/O operations as this more efficient use of the memory bandwidth as they can be coalesced into fewer operations as opposed to random ones.